This is a slope question.
Find x so that the line containing (1, 1) and (x, -2) will have a slope of 6.
my answer ended up being -2. but that didnt seem right.... help?

Answer by jim_thompson5910(35256)

m = (y2-y1)/(x2-x1)

6 = (-2-1)/(x-1)

6 = (-3)/(x-1)

6(x-1) = -3

6x-6 = -3

6x = -3+6

6x = 3

x = 3/6

x = 1/2